OpenStudy (anonymous):

Nesreen has a bottle of water that holds 2.4 L of water. She drinks half of it. How many mL remain? Explain your work.

OpenStudy (anonymous):

1/2 * 2,4 dm3 = 1,2 dm3 = 1,2 * 10^3 cm3 and cm3 = mL

OpenStudy (anonymous):

can you explain a little bit more easier for me

OpenStudy (anonymous):

ok so you have bottle with 2,4 L and L is equal to dm3, now you drink half of it so divide it with 2 and you will get 1,2 L or 1,2 dm3 now 1dm3 = 1*10^3 cm3 so you have 1,2 *10^3 cm3 or 1200 cm3 or 1200 mL

OpenStudy (anonymous):

oh!! thank you :)